# DiG-Net: Enhancing Quality of Life through Hyper-Range Dynamic Gesture Recognition in Assistive Robotics

### v1

@@ 2.1 Problem Formulation
[…]
Considering a sequence of past images enables the model to account for both the temporal and spatial dynamics of gestures, which is critical for differentiating between gestures that may appear similar when observed in a single frame.
2.2 Data Collection
@@
Figure 3: Example frames from the collected gesture dataset, showing different users, gestures, distances, and environments. The dataset includes both indoor and outdoor scenes with varying lighting and spatial arrangements to improve robustness and generalization.
2.3 Sequence Pre-processing
@@
A sequence $V_{t}$ contains $n$ images, some of which may be highly similar, leading to redundancy. Hence, the sequences in $\tilde{\mathcal{D}}$ are pre-processed to reduce the number of frames, resize them, and extract relevant features. Given a video $V_{t}$, the frames are reduced to $r<n$ representative frames using K-Means clustering. Each frame $\mathbf{I}_{j}\in V_{t}$ was processed through a pre-trained ResNet [59] to extract a low-dimensional feature vector $\mathbf{\phi}_{j}\in\mathbb{R}^{d}$. All feature vectors were clustered into $r$ clusters, and the frame corresponding to the feature vector closest to each cluster’s centroid was selected as the representative frame. In practice, a representative feature vector is chosen from cluster $C_{i}=\{\phi_{1},\phi_{2},\ldots\}$ according to
$\text{rep}(i)=\arg\min_{j}\|\mathbf{\phi}_{j}-\mathbf{C}_{i}\|_{2},$ (2)
@@
yielding $r$ representative frames $\{\mathbf{I}_{\text{rep},1},\mathbf{I}_{\text{rep},2},\ldots,\mathbf{I}_{\text{rep},r}\}$. YOLOv3 [60] was then used to detect the user within each frame and crop the background, enhancing focus, particularly when the user was far from the camera. Full-body detection was chosen, as it provides more stable and consistent localization at hyper-range distances. In contrast, detecting small body parts like hands results in very limited regions, making the input noisier and less informative due to the loss of fine details. While YOLOv3 assists in isolating the user, it occasionally fails at long distances due to resolution loss. This emphasizes the importance of the model’s robustness in handling degraded inputs for reliable recognition. To address cases where the bounding box did not fully capture the human body, the bounding box was extended while maintaining a constant aspect ratio. Specifically, the pixel extension added around the bounding box was $\frac{b}{a}$, where $b$ is the diagonal length of the bounding box, and $a$ is a predefined user-to-image ratio parameter. The resulting cropped image was resized to $224\times 224$ pixels to ensure uniformity across the dataset, maintaining a consistent size and aspect ratio.
@@
2.4 DiG-Net Model Framework
The proposed DiG-Net integrates DADA blocks with STG modules and Graph Transformer encoders to robustly handle spatial distortions and temporal dynamics of hand gestures at hyper-range distances. Individually, DADA blocks effectively compensate for physical attenuation and defocus blur, but cannot capture complex temporal dependencies. Conversely, STG modules and Graph Transformer encoders excel in modeling global spatio-temporal relationships but require well-aligned, noise-free inputs, making them sensitive to misalignment and distortions in low-resolution frames. DiG-Net addresses these complementary limitations by unifying depth-aware alignment, graph-based temporal reasoning, and global self-attention, thus enabling reliable recognition at distances of up to thirty meters. DiG-Net operates in a staged manner. Initially, input frames undergo successive DADA modules, which estimate motion-guided sampling offsets, perform ray-based feature warping, and correct for depth-related attenuation and defocus distortions. These depth-corrected features are then structured into a spatio-temporal graph, upon which the STG module executes spatial and temporal message passing to model local dynamics. Subsequently, Graph Transformer encoders apply multi-head self-attention across the graph nodes, effectively capturing long-range temporal dependencies and global contextual interactions. The self-attention mechanism of the Graph Transformer links early and late gesture phases, enhancing the importance of subtle deformation patterns identified through ray-sampling convolutions and resolving residual ambiguities inherent in distant, low-resolution inputs. Finally, the refined spatio-temporal features from the backbone network are concatenated into a tensor denoted by $\mathbf{Y}\in\mathbb{R}^{B\times c_{\mathrm{DADA}}\times T/4\times H/4\times W/4}$, where $B$ is the batch size, $c_{\mathrm{DADA}}$ represents the number of output channels after Graph Transformer processing, and $T$, $H$, and $W$ indicate the temporal length, spatial height, and spatial width of the input video sequence, respectively.
@@ 2.4 DiG-Net Model Framework
$S_{k}(x,y,\tau)=(x,y)+\frac{k}{z(x,y,\tau)}\,\hat{f},\quad\hat{f}=\frac{(u,v)}{\|(u,v)\|+\epsilon}$ (3)
@@
$F_{\mathrm{warp}}=\sum_{k=-K}^{K}g_{k}\bigl{(}z(x,y,\tau),\,\|(u,v)\|\bigr{)}\,\hat{\mathbf{F}}_{0}^{(k)}$ (4)
@@
$\mathbf{F}_{\mathrm{corr}}=F_{\mathrm{warp}}\,\exp\!\bigl{(}\eta\,z(x,y,\tau)\bigr{)}$ (5)
with $\eta$ the Beer–Lambert attenuation coefficient [61]. After spatio‐temporal graph and Graph Transformer processing, the depth‐corrected features are aggregated over space and time:
$\mathbf{h}_{i}=\frac{1}{T^{\prime}H^{\prime}W^{\prime}}\sum_{\tau=1}^{T^{\prime}}\sum_{x=1}^{H^{\prime}}\sum_{y=1}^{W^{\prime}}\mathbf{F}_{\mathrm{corr}}^{(i)}(x,y,\tau),$ (6)
@@
$\mathcal{M}(\rho_{i},\xi_{i})=\gamma_{1}(1-e^{-\mu\rho_{i}})+\gamma_{2}Q+\gamma_{3}\left(1-e^{-\lambda\xi_{i}}\right),$ (7)
where Q = $\left(1-\frac{1}{1+(\rho_{i}/\rho_{0})^{2}}\right)$ and $\rho_{i}$ is the distance of the gesture, $\xi_{i}$ is the average motion magnitude from the optical flow, and $\mu,\rho_{0},\lambda$ are learnable attenuation parameters. The overall RSTDAL loss is then given by:
$\mathcal{L}_{\text{RSTDAL}}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{G}{G+\sum_{j\neq y_{i}}\exp\left(s\langle\mathbf{e}_{i},\boldsymbol{\theta}_{j}\rangle\right)},$ (8)
@@
3 Model Evaluation
In this section, we evaluate the proposed DiG-Net framework for recognizing dynamic gestures used to naturally guide a robot. All computations were conducted on a Linux Ubuntu 18.04 LTS system equipped with an Intel Xeon Gold 6230R CPU (20 cores at 2.1 GHz) and four NVIDIA GeForce RTX 2080TI GPUs, each with 11 GB of RAM. Hyperparameter tuning was performed using Ray-Tune [62] to all models.
3.1 Gestures
The evaluation focuses on $m=13$ distinct gesture classes. Of the 13, eight are dynamic, seen in Figure 4, and include: go-back with a forward and backward motion of an open hand, palm facing outward; go-up with an upward motion of an open hand, palm facing upward; go-down with a downward motion of an open hand, palm facing downward; move-right with a horizontal sweeping motion of the open hand, palm facing right; move-left with horizontal sweeping motion of the open hand, palm facing left; turn-around with a circular motion of the index finger, pointing upwards; beckoning where the palm is facing upward and the fingers are rhythmically flexed and extended; follow-me where the open palm is repeatedly tapped on the head of the user.
@@ 3.1 Gestures
Figure 4: The eight dynamic gestures used in the analysis include: (a) beckoning, (b) go-back, (c) move-right, (d) move-left, (e) turn-around, (f) follow-me, (g) go-down and (h) go-up.
3.2 Dataset
@@
Dataset $\mathcal{D}$ was collected by recording video samples of the $m=13$ gestures at a distance range of $\rho\in[2,30]$. Gestures were exhibited in diverse environments, including both indoor and outdoor settings, varying lighting, and diverse backgrounds. In addition, 16 different participants contributed gesture data to ensure variability in gesture performance. Each participant performed each gesture multiple times at different measured distances. The data collection process involved using a standard RGB camera with a resolution of $640\times 480$ pixels taken at $21$ frames per second. Each gesture was manually annotated according to its class $o_{i}$ and the distance $\rho_{i}$ at which it was performed, resulting in labeled data suitable for supervised learning. This annotation allowed for a detailed analysis of model performance across varying distances, as well as the development of distance-specific recognition strategies. The collected dataset $\mathcal{D}$ yielded $N=3,240$ video samples of hand gestures, each lasting 4 seconds with up to $n=84$ frames. After feature extraction as described in Section 2.3, the number of frames was reduced to $k=8$ per video. With further augmentation, dataset $\tilde{\mathcal{D}}$ includes $M=4,790$ samples. An additional test set of $K=458$ labeled and processed videos was recorded in distinct environments to evaluate the model’s performance.
3.3 Comparative Evaluation
@@
We now evaluate the proposed DiG-Net model compared to existing state-of-the-art gesture recognition frameworks. The DiG-Net model was trained with the Lion optimizer [63], chosen for its efficiency and convergence stability, using an initial learning rate of 0.0037. The learning rate was gradually reduced using a cosine annealing schedule to ensure stable convergence. The model was trained for 100 epochs with a batch size of $B=16$, using the proposed RSTDAL loss function (8), which adaptively adjusts the angular margin based on the gesture distance $\rho_{i}\in[2,30]$ and motion magnitude $\xi_{i}$ through the learnable function $\mathcal{M}(\rho_{i},\xi_{i})$. The margin function incorporated fixed attenuation parameters of $\mu=0.1$ and $\lambda=0.2$, with a reference distance $\rho_{0}=16$ used in the non-linear margin term to center the response. The weights of the three margin components were set to $\gamma_{1}=0.4$, $\gamma_{2}=0.5$, and $\gamma_{3}=0.2$, respectively, to balance the influence of exponential decay over distance, the non-linear reference curve, and motion magnitude. These hyperparameters were optimized using RayTune [64] to balance the emphasis on long-range samples and ensure stable training. To prevent overfitting, dropout was applied at the output of the Graph Transformer encoder, and L2 regularization was used during training. Early stopping was employed by monitoring the validation loss to determine the optimal stopping epoch. The DiG-Net model is compared to the: Swin Transformer [65] which uses shifted windows for efficient spatial feature extraction; ViViT [66], a Transformer for video analysis capturing spatial and temporal relationships; TimeSformer [67] which separates temporal and spatial attention; MViT [68], a multiscale Transformer for spatiotemporal resolutions; I3D [69], a 3D convolutional network for spatiotemporal learning; and, X3D [70], an efficient 3D convolutional model balancing speed and accuracy.
@@
$\text{DWA}=\frac{1}{K}\sum_{i=1}^{K}\mathbb{I}(\tilde{o}_{i}=o_{i})w_{i},$ (9)
@@
$\text{GSS}=\frac{1}{K}\sum_{i=1}^{K}\left(\frac{1}{n_{i}}\sum_{j=n}^{n_{i}}\mathbb{I}(\tilde{o}_{i,j}=o_{i})\right),$ (10)
@@
Table 3: Evaluation results for different dynamic gesture recognition models
Model Success rate (%) DWA GSS $F_{1}$ score mAP (%) Inference Time (ms)
Swin [65] 80.5 0.84 0.85 0.83 78.1 38
ViViT [66] 78.3 0.82 0.84 0.80 77.5 44
TimeSformer [67] 83.4 0.85 0.87 0.85 81.3 49
MViT [68] 87.9 0.88 0.90 0.89 85.1 26
I3D [69] 84.3 0.86 0.88 0.85 82.4 31
X3D [70] 86.2 0.87 0.89 0.87 78.8 29
CorrNet [71] 82.1 0.84 0.88 0.83 79.3 45
GestFormer [72] 85.4 0.83 0.89 0.85 81.8 27
DiG-Net 97.3 0.92 0.96 0.93 94.9 35
3.4 DiG-Net analysis
We further analyze the performance of the proposed DiG-Net model. Figure 5 presents the gesture recognition success rate with respect to the distance $\rho$ between the user and the camera. The success rate gradually decreases but remains relatively high in the desired range of up to 30 meters. This demonstrates the robustness of the DiG-Net model in recognizing gestures at hyper-range, though performance diminishes as distance increases due to factors such as reduced resolution and increased visual noise. Figure 6 presents the confusion matrix for the DiG-Net model over the test data and for all 13 gesture classes. While some gestures may be visually similar, the temporal context modeling enabled by the Graph Transformer encoder and spatio-temporal graph structure helps disambiguate subtle motion patterns, reducing the likelihood of misclassification.
Figure 5: Gesture recognition success rate of the DiG-Net model with regard to the distance $\rho$ of the user from the camera.
Figure 6: Confusion matrix for the gesture classification with the DiG-Net model across 13 gesture classes.

### v2

@@ 2.1. Problem Formulation
(1) $j^{*}=\underset{j}{\mathrm{argmax}}\;P(O_{j}\mid V_{t}),\quad j=1,\dots,m~.$
Considering a sequence of past images enables the model to account for both the temporal and spatial dynamics of gestures, which is critical for differentiating between gestures that may appear similar when observed in a single frame. Throughout this paper we consider a fixed vocabulary of $m=13$ gesture classes comprising eight dynamic gestures {go-back, go-up, go-down, move-right, move-left, turn-around, beckoning, follow-me}, four static gestures {pointing, thumbs-up, thumbs-down, stop}, and a null class indicating no intentional gesture. This vocabulary is used consistently for data collection, training, and evaluation; the dynamic gestures are illustrated in Figure 4, with detailed definitions provided later in Section 3.1.
2.2. Data Collection
@@
Figure 3. Example frames from the collected gesture dataset showing different users, gestures, and distances in indoor and outdoor environments.
Sample frames from the DiG-Net gesture dataset illustrating multiple users performing various dynamic gestures across different distances, lighting conditions, and backgrounds. The dataset includes both indoor and outdoor scenes with diverse spatial arrangements to enhance robustness and generalization for hyper-range gesture recognition.
2.3. User Study on Human Recognition of Long-Range Gestures
To better understand how humans perceive gestures at extended distances, we conducted a user study that complements the technical evaluation of DiG-Net by examining the perceptual challenges humans face under similar conditions. The goal was to assess human recognition performance and subjective confidence when identifying static and dynamic gestures at varying distances, thereby providing a behavioral reference for interpreting the model’s capabilities. A total of ten participants (three female, seven male) took part in the experiment. The study followed a dense repeated-measures design: each participant evaluated approximately 50 randomly sampled trials spanning all gesture classes and the three distance conditions, yielding $\sim$500 individual human observations. This supports robust within-subject trend estimation across distance, while we note that N=10N{=}10 limits population-level generalizability; larger and more diverse cohorts are left for future work.
Each participant sat in front of a computer and was presented with both single-frame images and short video clips depicting human gestures recorded at three distances—short (5–10 m), medium (15–20 m), and long (25–30 m). All participants completed the study under controlled viewing conditions, using a full-HD monitor in a quiet, evenly lit indoor environment with minimal visual distractions. The gestures corresponded to those used in the DiG-Net dataset, including ”stop”, ”come”, ”go back”, ”turn left”, and ”turn right”. For each presented sample, participants were asked to (1) identify which gesture was shown, (2) press a key as soon as they recognized it (reaction time), and (3) rate their confidence in the identification on a 7-point Likert scale. The experiment revealed a clear distance-dependent effect on recognition accuracy and confidence. Participants achieved near-perfect recognition for both static and dynamic gestures at short range, while performance degraded substantially beyond 20 m, particularly for static frames where spatial details were lost. In contrast, dynamic gesture clips maintained relatively high recognition rates even at long distances, suggesting that motion cues play a key role in preserving interpretability. Qualitative feedback indicated that participants perceived dynamic gestures as ”more natural and expressive”, and several noted that motion helped them infer the actor’s intention even when fine-grained shape information was unclear. Confusion matrices further showed that errors primarily occurred between visually similar gestures, such as ”stop” and ”go back”, especially at long range. These findings provide human-perception grounding for DiG-Net’s spatio-temporal design: while humans struggle with static, distance-degraded inputs, temporal motion dynamics enable more reliable interpretation. This supports the model’s emphasis on depth-compensated and motion-aware representations. Moreover, the results underscore the broader significance of long-range gestures as a natural, intuitive communication channel for guiding robots and establishing mutual understanding in assistive and collaborative scenarios. A quantitative summary of user study outcomes is provided in Table 5 within the Evaluation section.
2.4. Sequence Pre-processing
@@
A sequence $V_{t}$ contains $n$ images, some of which may be highly similar, leading to redundancy. Hence, the sequences in $\tilde{\mathcal{D}}$ are pre-processed to reduce the number of frames, resize them, and extract relevant features. Given a video $V_{t}$, the frames are reduced to $r<n$ representative frames using K-Means clustering. Each frame $\mathbf{I}_{j}\in V_{t}$ was processed through a pre-trained ResNet (He et al., 2016) to extract a low-dimensional feature vector $\mathbf{\phi}_{j}\in\mathbb{R}^{d}$. All feature vectors were clustered into $r$ clusters, and the frame corresponding to the feature vector closest to each cluster’s centroid was selected as the representative frame. In practice, a representative feature vector is chosen from cluster $C_{i}=\{\phi_{1},\phi_{2},\ldots\}$ according to
(2) $\text{rep}(i)=\arg\min_{j}\|\mathbf{\phi}_{j}-\mathbf{C}_{i}\|_{2},$
@@
yielding $r$ representative frames $\{\mathbf{I}_{\text{rep},1},\mathbf{I}_{\text{rep},2},\ldots,\mathbf{I}_{\text{rep},r}\}$. YOLOv3 (Redmon and Farhadi, 2018) was then used to detect the user within each frame and crop the background, enhancing focus, particularly when the user was far from the camera. Full-body detection was chosen, as it provides more stable and consistent localization at hyper-range distances. In contrast, detecting small body parts like hands results in very limited regions, making the input noisier and less informative due to the loss of fine details. While YOLOv3 assists in isolating the user, it occasionally fails at long distances due to resolution loss. This emphasizes the importance of the model’s robustness in handling degraded inputs for reliable recognition. To address cases where the bounding box did not fully capture the human body, the bounding box was extended while maintaining a constant aspect ratio. Specifically, the pixel extension added around the bounding box was $\frac{b}{a}$, where $b$ is the diagonal length of the bounding box, and $a$ is a predefined user-to-image ratio parameter. The resulting cropped image was resized to $224\times 224$ pixels to ensure uniformity across the dataset, maintaining a consistent size and aspect ratio.
@@
2.5. DiG-Net Model Framework
The proposed DiG-Net integrates DADA blocks with STG modules and Graph Transformer encoders to robustly handle spatial distortions and temporal dynamics of hand gestures at hyper-range distances. Our DADA modules build upon the deformable convolution framework introduced by Dai et al. (Dai et al., 2017), extending it by conditioning the learned offsets on estimated depth and motion cues. This depth-aware adaptation enables the model to handle better spatial distortions and attenuation effects unique to hyper-range gesture recognition. Individually, DADA blocks effectively compensate for physical attenuation and defocus blur, but cannot capture complex temporal dependencies. Conversely, STG modules and Graph Transformer encoders excel in modeling global spatio-temporal relationships but require well-aligned, noise-free inputs, making them sensitive to misalignment and distortions in low-resolution frames. DiG-Net addresses these complementary limitations by unifying depth-aware alignment, graph-based temporal reasoning, and global self-attention, thus enabling reliable recognition at distances of up to thirty meters. DiG-Net operates in a staged manner. Initially, input frames undergo successive DADA modules, which estimate motion-guided sampling offsets, perform ray-based feature warping, and correct for depth-related attenuation and defocus distortions. These depth-corrected features are then structured into a spatio-temporal graph, upon which the STG module executes spatial and temporal message passing to model local dynamics. Subsequently, Graph Transformer encoders apply multi-head self-attention across the graph nodes, effectively capturing long-range temporal dependencies and global contextual interactions. The self-attention mechanism of the Graph Transformer links early and late gesture phases, enhancing the importance of subtle deformation patterns identified through ray-sampling convolutions and resolving residual ambiguities inherent in distant, low-resolution inputs. Finally, the refined spatio-temporal features from the backbone network are concatenated into a tensor denoted by $\mathbf{Y}\in\mathbb{R}^{B\times c_{\mathrm{DADA}}\times T/4\times H/4\times W/4}$, where $B$ is the batch size, $c_{\mathrm{DADA}}$ represents the number of output channels after Graph Transformer processing, and $T$, $H$, and $W$ indicate the temporal length, spatial height, and spatial width of the input video sequence, respectively.
@@ 2.5. DiG-Net Model Framework
(3) $S_{k}(x,y,\tau)=(x,y)+\frac{k}{z(x,y,\tau)}\,\hat{f},\quad\hat{f}=\frac{(u,v)}{\|(u,v)\|+\epsilon}$
@@
(4) $F_{\mathrm{warp}}=\sum_{k=-K}^{K}g_{k}\bigl(z(x,y,\tau),\,\|(u,v)\|\bigr)\,\hat{\mathbf{F}}_{0}^{(k)}$
@@
(5) $\mathbf{F}_{\mathrm{corr}}=F_{\mathrm{warp}}\,\exp\!\bigl(\eta\,z(x,y,\tau)\bigr)$
with $\eta$ the Beer–Lambert attenuation coefficient (Swinehart, 1962). After spatio‐temporal graph and Graph Transformer processing, the depth‐corrected features are aggregated over space and time:
(6) $\mathbf{h}_{i}=\frac{1}{T^{\prime}H^{\prime}W^{\prime}}\sum_{\tau=1}^{T^{\prime}}\sum_{x=1}^{H^{\prime}}\sum_{y=1}^{W^{\prime}}\mathbf{F}_{\mathrm{corr}}^{(i)}(x,y,\tau),$
@@
(7) $\mathcal{M}(\rho_{i},\xi_{i})=\gamma_{1}(1-e^{-\mu\rho_{i}})+\gamma_{2}Q+\gamma_{3}\left(1-e^{-\lambda\xi_{i}}\right),$
where Q = $\left(1-\frac{1}{1+(\rho_{i}/\rho_{0})^{2}}\right)$ and $\rho_{i}$ is the distance of the gesture, $\xi_{i}$ is the average motion magnitude from the optical flow, and $\mu,\rho_{0},\lambda$ are learnable attenuation parameters. The intuition behind RSTDAL is to embed a physical prior directly into the decision geometry of the classifier. In hyper-range scenarios, signal quality degrades monotonically with distance, naturally reducing the separability between gesture classes. Standard losses often fail to prioritize these ”hard” examples. RSTDAL addresses this by utilizing the Beer–Lambert attenuation term not to restore the signal, but to dynamically adjust the classification margin. By enforcing a larger margin for samples affected by high attenuation (distance) or low motion intensity, the objective function penalizes ambiguous embeddings in difficult conditions. This drives the network to learn distance-robust representations that rely on stable spatio-temporal dynamics rather than fine-grained features that are susceptible to degradation. The overall RSTDAL loss is then given by:
(8) $\mathcal{L}_{\text{RSTDAL}}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{G}{G+\sum_{j\neq y_{i}}\exp\left(s\langle\mathbf{e}_{i},\boldsymbol{\theta}_{j}\rangle\right)},$
@@
3. Model Evaluation
In this section, we evaluate the proposed DiG-Net framework for recognizing dynamic gestures used to naturally guide a robot. All computations were conducted on a Linux Ubuntu 18.04 LTS system equipped with an Intel Xeon Gold 6230R CPU (20 cores at 2.1 GHz) and four NVIDIA GeForce RTX 2080TI GPUs, each with 11 GB of RAM. Hyperparameter tuning was performed using Ray-Tune (Liaw et al., 2018) for all models.
3.1. Gestures
The evaluation focuses on $m=13$ distinct gesture classes. Of the 13, eight are dynamic, illustrated in Figure 4, and include: go-back with a forward and backward motion of an open hand, palm facing outward; go-up with an upward motion of an open hand, palm facing upward; go-down with a downward motion of an open hand, palm facing downward; move-right with a horizontal sweeping motion of the open hand, palm facing right; move-left with horizontal sweeping motion of the open hand, palm facing left; turn-around with a circular motion of the index finger, pointing upwards; beckoning where the palm is facing upward and the fingers are rhythmically flexed and extended; follow-me where the open palm is repeatedly tapped on the head of the user.
@@ 3.1. Gestures
Figure 4. The eight dynamic gestures used in the analysis: (a) beckoning, (b) go-back, (c) move-right, (d) move-left, (e) turn-around, (f) follow-me, (g) go-down, and (h) go-up.
A visual illustration of the eight dynamic hand gestures in the DiG-Net dataset. Each subimage shows a person performing one of the gestures: beckoning, go-back, move-right, move-left, turn-around, follow-me, go-down, and go-up. The gestures are depicted as sequences of hand and arm movements, providing visual references for the actions analyzed in the study.
3.2. Dataset
@@
Dataset $\mathcal{D}$ was collected by recording video samples of the $m=13$ gestures at a distance range of $\rho\in[2,30]$. Gestures were exhibited in diverse environments, including both indoor and outdoor settings, varying lighting, and diverse backgrounds. In addition, 16 different participants (11 males and 5 females, aged 25–44 years) contributed gesture data to ensure variability in gesture performance. Each participant performed each gesture multiple times at different measured distances. The data collection process involved using a standard RGB camera with a resolution of $640\times 480$ pixels taken at $21$ frames per second. Each gesture was manually annotated according to its class $o_{i}$ and the distance $\rho_{i}$ at which it was performed, resulting in labeled data suitable for supervised learning. This annotation allowed for a detailed analysis of model performance across varying distances, as well as the development of distance-specific recognition strategies. The collected dataset $\mathcal{D}$ yielded $N=3,240$ video samples of hand gestures, each lasting 4 seconds with up to $n=84$ frames. After feature extraction as described in Section 2.4, the number of frames was reduced to $k=8$ per video. With further augmentation, dataset $\tilde{\mathcal{D}}$ includes $M=4,790$ samples. An additional test set of $K=458$ labeled and processed videos was recorded in distinct environments to evaluate the model’s performance.
3.3. Comparative Evaluation
@@
We now evaluate the proposed DiG-Net model compared to existing state-of-the-art gesture recognition frameworks. The DiG-Net model was trained with the Lion optimizer (Chen et al., 2023), chosen for its efficiency and convergence stability, using an initial learning rate of 0.0037. The learning rate was gradually reduced using a cosine annealing schedule to ensure stable convergence. The model was trained for 100 epochs with a batch size of $B=16$, using the proposed RSTDAL loss function (8), which adaptively adjusts the angular margin based on the gesture distance $\rho_{i}\in[2,30]$ and motion magnitude $\xi_{i}$ through the parameterized margin function $\mathcal{M}(\rho_{i},\xi_{i})$. The margin function used $\mu=0.1$ and $\lambda=0.2$ to control the saturation rates of the distance-driven and motion-driven margin terms, respectively, with a reference distance $\rho_{0}=16$ used in the non-linear margin term to center the response. The weights of the three margin components were set to $\gamma_{1}=0.4$, $\gamma_{2}=0.5$, and $\gamma_{3}=0.2$, respectively, to balance the influence of exponential decay over distance, the non-linear reference curve, and motion magnitude. These hyperparameters were optimized using RayTune (Liaw et al., 2018) to balance the emphasis on long-range samples and ensure stable training.
We analyzed the sensitivity of the RSTDAL hyperparameters. Specifically, $\mu$ and $\lambda$ control the saturation rate of the distance- and motion-driven margins. High $\lambda$ values cause the motion term $(1-e^{-\lambda\xi_{i}})$ to saturate rapidly, potentially over-penalizing subtle gestures, while low values render the motion cue negligible. Similarly, an excessively high $\mu$ causes the distance term $(1-e^{-\mu\rho_{i}})$ to dominate the gradient and bias optimization toward hyper-range samples, whereas a low $\mu$ fails to enforce sufficient separability at range. Additionally, $\rho_{0}$ sets the reference distance scale for the term $Q$, and $\gamma_{1,2,3}$ balances the contribution of the three margin components. The final values were selected via RayTune validation to balance validation performance across the full 3–30m range.
To prevent overfitting, dropout was applied at the output of the Graph Transformer encoder, and L2 regularization was used during training. Early stopping was employed by monitoring the validation loss to determine the optimal stopping epoch. The DiG-Net model is compared to the: Swin Transformer (Liu et al., 2022) which uses shifted windows for efficient spatial feature extraction; ViViT (Arnab et al., 2021), a Transformer for video analysis capturing spatial and temporal relationships; TimeSformer (Bertasius et al., 2021) which separates temporal and spatial attention; MViT (Fan et al., 2021), a multiscale Transformer for spatiotemporal resolutions; I3D (Carreira and Zisserman, 2017), a 3D convolutional network for spatiotemporal learning; and, X3D (Feichtenhofer, 2020), an efficient 3D convolutional model balancing speed and accuracy.
@@
(9) $\text{DWA}=\frac{1}{K}\sum_{i=1}^{K}\mathbb{I}(\tilde{o}_{i}=o_{i})w_{i},$
@@
(10) $\text{GSS}=\frac{1}{K}\sum_{i=1}^{K}\left(\frac{1}{n_{i}}\sum_{j=n}^{n_{i}}\mathbb{I}(\tilde{o}_{i,j}=o_{i})\right),$
@@
Table 3. Evaluation results for different dynamic gesture recognition models
[…]
3.4. DiG-Net analysis
We further analyze the performance of the proposed DiG-Net model. Figure 5 illustrates the gesture recognition success rate with respect to the distance $\rho$ between the user and the camera. The success rate gradually decreases but remains relatively high in the desired range of up to 30 meters. This demonstrates the robustness of the DiG-Net model in recognizing gestures at hyper-range, though performance diminishes as distance increases due to factors such as reduced resolution and increased visual noise. Figure 6 presents the confusion matrix for the DiG-Net model over the test data and for all 13 gesture classes. While some gestures may be visually similar, the temporal context modeling enabled by the Graph Transformer encoder and spatio-temporal graph structure helps disambiguate subtle motion patterns, reducing the likelihood of misclassification.
Figure 5. Gesture recognition success rate of the DiG-Net model as a function of the user’s distance ($\rho$) from the camera. Performance gradually decreases at longer ranges due to lower image resolution and atmospheric effects.
A line plot showing the recognition success rate of the DiG-Net model versus the user’s distance from the camera. Accuracy remains high at close and mid ranges and slowly declines as distance increases, demonstrating the model’s robustness up to approximately 30 meters. Axes are labeled with distance on the x-axis and success rate on the y-axis.
Figure 6. Confusion matrix for the gesture classification with the DiG-Net model across 13 gesture classes.
A heatmap representing the confusion matrix for 13 dynamic gesture classes using the DiG-Net model. The x-axis represents predicted gesture classes, and the y-axis represents true classes. Most values are concentrated along the diagonal, indicating strong classification accuracy, with minimal off-diagonal errors showing low confusion between gestures.
To complement the quantitative evaluation, a qualitative analysis was conducted to better illustrate DiG-Net’s behavior under varying visual conditions. The confusion patterns in Figure 6, together with the robustness results in Tables 69, reveal consistent trends in both success and failure regimes. DiG-Net demonstrates the highest reliability under hyper-range conditions, where fine-grained appearance cues vanish, and temporal information becomes the dominant discriminative factor. In these scenarios, the model maintains clear separability between gestures that are visually similar but differ in motion dynamics, such as go-back and stop, or turn-around and pointing. This behavior reflects the contribution of the depth-aware alignment and spatio-temporal reasoning modules, which jointly stabilize degraded inputs and preserve motion coherence. Failure cases primarily occur under extreme environmental degradation, consistent with the performance decline observed under severe clutter/interference and strong optical degradation conditions (Tables 7 and 9). In such situations, low signal-to-noise ratios obscure fine motion details, occasionally causing confusion between gestures with similar overall trajectories when hand visibility is reduced. Overall, these qualitative findings align with the quantitative robustness analysis and highlight both the strengths and the remaining limitations of DiG-Net in real-world long-range conditions.